OpenStudy (anonymous):

Check my work please? What is the solution to the rational equation x over 5 minus 2 over 5 = 7 ? x = 31 x = 33 x = 37 x = 39 I got B..

jimthompson5910 (jim_thompson5910):

B is incorrect

OpenStudy (anonymous):

ughwhgohr. can you explain?

OpenStudy (campbell_st):

not quite right the question can be written as \[\frac{x - 2}{5} = 7\]

jimthompson5910 (jim_thompson5910):

campbell_st has the right idea and you can multiply both sides by 5 to get x-2 = 7*5 x-2 = 35 x = ??

OpenStudy (anonymous):

they are different divisions. x-5/( -2/5)=7

OpenStudy (anonymous):

37?

jimthompson5910 (jim_thompson5910):

37 is correct

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw